OpenStudy (candycane):

if someone could help me figure out what formula to use for the following problem thatd be awesome Are the two triangles similar how do you know?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

not sure if you are looking for a "formula" check to see if the ratios are the same

OpenStudy (candycane):



OpenStudy (misty1212):

what is the missing angle from each triangle/ hint: the three angles in each have to add up to 180

OpenStudy (misty1212):

also ignore what i said about the ratios, i thought you had the sides, but you have the angles check to see if all angles are the same

OpenStudy (misty1212):



OpenStudy (candycane):

if i had53 and 60 I get 113 180-113= 60 60+67=127 180-127 is 50, the angles seem to matchup so theyre similar?

OpenStudy (misty1212):

yes, all angles are the same the missing angle from the first one is \(180-60-53=67\) and the missing angle from the second is \(180-60-67=53\)

OpenStudy (misty1212):

and yes, that means they are similar

OpenStudy (candycane):

Awesome thank you, I must have been over thinking it, as usual haha than kyou so much :)

OpenStudy (misty1212):

https://www.youtube.com/watch?v=XPgGjUSEWss \[\color\magenta\heartsuit\]